for a zero order reaction , K= 1 *10 power -3 mol/lit/sec.If initial conc of the reactant is 1 mol/lit,the conc after 10 min would be.?

Grade:12

3 Answers

Naveen Kumar
9 years ago
For zero order reaction, -d[A]/dt=k* [A]0=k, integrating the reaction and putting values of [A] at t=0 and t=10min, we would get the relation,
[A]0-[A]t=k*t
where[A]0= concentration at t=0
and[A]t=concentration at t(=10 min)
k=rate constant
putting the given values, 1mol/L – [A]10min = 10-3 *(10*60) mol/L
So, concentration after 10 min[A]10min =1-600/1000=1-0.6 = 0.4 mol/L
Shriya Mehrotra
9 years ago
thanks sir.
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

For zero order reaction,
-d[A]/dt=k* [A]0=k,
integrating the reaction and putting values of [A] at t=0 and t=10min,
we would get the relation,
[A]0-[A]t=k*t where
[A]0= concentration at t=0 and
[A]t=concentration at t(=10 min) k=rate constant putting the given values,
1mol/L – [A]10min = 10-3 *(10*60) mol/L
So, concentration after 10 min[A]10min =1-600/1000=1-0.6 = 0.4 mol/L

Thanks and Regards